# The impact of $\gamma N$ and $\gamma^{*}N$ interactions on our understanding of nucleon excitations

### v1

@@ 4.6   Baryons in lattice QCD
Lattice QCD is a nonperturbative ab-initio method because in principle, apart from supplying the current-quark masses and setting a scale, the only information needed to compute baryon observables is the QCD action [801, 802]. The basic idea in lattice QFT is to discretize Euclidean spacetime and consider a four-dimensional box with lattice spacing $a$, length $L$ and appropriate boundary conditions. Correlation functions inside that box are then calculated through the Euclidean version of the path integral in Eq. (4.7). The path integral over the quark and antiquark fields can be rewritten as a fermion determinant, so that the remaining integral amounts to statistically sampling over many gauge (i.e., gluon) field configurations. The typical workflow of lattice calculations consists of three steps: generating the gauge-field ensembles that enter in the path integral; ‘measuring’ the correlation functions $\langle\mathcal{O}\rangle$ on those ensembles; and extracting physical observables from the correlation functions. The last part comes with several complications mainly induced by the finite volume ($L<\infty$), the non-zero lattice spacing ($a>0$), and the difficulties in going to physical pion masses due to the accumulating statistical noise.
The starting point for extracting the baryon spectrum on the lattice is again Fig. 4.3. To begin with, suppose we start with some (gauge-dependent) correlation function that permits baryon poles, like the $qqq$ correlation function in the second row on the right. By contracting it with appropriate structures in Dirac, color and flavor space, we arrive at a (gauge-invariant) two-point correlator which is shown in the third row. This two-point function only depends on one variable $z=x-y$ and can be interpreted as a baryon-baryon correlator $G(z)=\langle\mathcal{O}_{i}(x)\,\bar{\mathcal{O}}_{j}(y)\rangle$, where $\mathcal{O}_{i}(x)$ is some gauge-invariant combination of three quark operators. In Euclidean momentum space, this correlator depends on the total baryon momentum $P$ and inherits the poles from the original correlation function at $P^{2}=-m_{\lambda}^{2}$:
@@
where $E_{\lambda}^{2}={\mbox{\boldmath$P$}}^{2}+m_{\lambda}^{2}$ and the overlap $R_{\lambda}$ is the product of the contracted Bethe-Salpeter wave functions. When taking a Fourier transform in $P_{4}$, a pole in momentum space turns into an exponential decay in Euclidean time $z_{4}=\tau$:
@@
$G(\tau,{\mbox{\boldmath$P$}})\stackrel{{\scriptstyle P^{2}\to-m_{\lambda}^{2}}}{{\relbar\joinrel\relbar\joinrel\relbar\joinrel\relbar\joinrel\relbar\joinrel\relbar\joinrel\rightarrow}}R_{\lambda}\int\frac{dP_{4}}{2\pi}\frac{e^{iP_{4}z_{4}}}{P_{4}^{2}+E_{\lambda}^{2}}=\frac{R_{\lambda}}{2E_{\lambda}}e^{-E_{\lambda}\tau}\,.$ (4.37)
[…]
@@
While the above equations sketch the basic ideas, a clean extraction of hadronic states from lattice QCD requires overcoming significant technical challenges. Over the years, various methods have been developed to address these issues and reduce the computational cost, including distillation [803, 804], noise reduction techniques [805], the inclusion of large operator bases also with multihadron operators to solve generalized eigenvalue problems, and matching the irreducible representations of the cubic group to physical $J^{P}$ quantum numbers; see e.g. [806, 807, 808, 809] for details. From the perspective of Fig. 4.3, the inclusion of multihadron and gluonic operators for solving generalized eigenvalue problems of two-point functions $\langle\mathcal{O}_{i}\,\bar{\mathcal{O}}_{j}\rangle$ amounts to finding a given hadron pole in a matrix of $qqq$, $(qqq)(q\bar{q})$, $(qqq)g$correlation functions, which therefore also provides insight into their multiquark and gluonic admixtures. In this way, for instance, the existence of hybrid baryons with significant gluonic components has been demonstrated [810].
@@
Lattice calculations of stable hadrons are well established by now, see e.g. [795, 796, 797, 811, 812, 813, 798, 799]. In the absence of electroweak interactions, the ground-state octet baryons $N$, $\Lambda$, $\Sigma$, $\Xi$ and the decuplet $\Omega$ baryon are stable under the strong interaction. Thanks to improved algorithms, calculations of their masses have also become possible at physical pion masses as shown in the left panel of Fig. 4.15. This also allows one to investigate precision effects like isospin breaking and electromagnetic corrections to the masses [814, 815, 816, 817]. The FLAG (Flavour Lattice Averaging Group) review [818, 819] provides a comprehensive meta-analysis of results for several hadronic quantities ensuring strict quality criteria such as continuum extrapolations, finite-volume corrections and renormalization procedures.
@@
Figure 4.16: (color online) Left and center: Sketch of the Lüscher method. From the knowledge of the finite-volume spectrum at different lattice volumes one can reconstruct the phase shifts and hence the scattering amplitude. Right: Sketch of the $N\pi$, $N\pi\pi$, …thresholds (dashed) plotted over the squared pion mass. The squares are the experimental values for the nucleon, $\Delta$ and Roper mass at the physical pion mass. The solid curves are parametrizations of lattice results for $m_{N}$ and $m_{\Delta}$ (see [692] for references).
@@
In contrast to stable ground states, the extraction of unstable hadrons above thresholds is substantially more difficult on the lattice. The reason for this is the finite lattice volume $L$, which does not admit branch cuts, Riemann sheets or resonance poles in scattering amplitudes [820]. Instead, the energy levels calculated on the lattice correspond to discrete poles on the real axis, which are the finite-volume analogues of the branch cuts in the infinite volume, i.e., the multiparticle continuum. As such, the energy levels above thresholds also cannot be directly interpreted as physical states.
@@
One way to deal with this is the Lüscher formalism [652, 653] including its various generalizations [821, 822, 823, 824, 825, 826, 827, 828, 829], see [802, 830, 831] for reviews. The basic idea is best explained with the help of a Bethe-Salpeter equation. Consider an elastic $2\to 2$ scattering amplitude such as $N\pi\to N\pi$ (or a coupled system of such scattering amplitudes), which creates intermediate poles for the nucleon and its excitations. Like in Eq. (4.15), we write down the generic BSE for this amplitude, once in the infinite volume and once in a finite volume (the latter with subscript $L$):
@@
Over the past decade, large progress has been made in lattice QCD calculations of hadron masses above two-particle thresholds; see [802, 832, 833, 683, 834, 604] for overviews. Most practical applications of the Luescher method to date have focused on the meson sector. The case of elastic scattering applies to resonances with a single decay channel, such as the $\rho$ and $K^{\ast}$ which decay almost exclusively into $\pi\pi$ and $K\pi$, respectively. Other resonances like the $\sigma$ and $\kappa$ mesons [835, 836, 837] and a number of states in the heavy-meson spectrum, including exotic candidates [838, 839, 840, 841, 842, 843, 844, 845, 846, 847], have also been studied by various lattice collaborations. Since most resonances are inelastic and couple to multiple decay channels, also coupled-channel generalizations of the Lüscher formalism to inelastic scattering have been applied to a variety of systems [848, 849, 850, 851, 852, 853, 854, 855]. These investigations enable the determination of resonance pole positions in the complex plane and provide insight into their dependence on the current-quark masses.
@@
The difficulty is that each additional threshold complicates the situation substantially. For unstable states above three-particle thresholds, the corresponding BSE turns into a genuine three-body problem. The generalization of the Lüscher condition to states above three-particle thresholds has been worked out by several groups [856, 857, 858, 859, 860, 861, 862, 863, 864, 865], see [830, 831] for reviews. This has found various recent applications in the meson sector, e.g., for $\pi\pi\pi$ and $KKK$ systems [866, 867, 868, 869, 870], from where one can determine the pole positions of the $a_{1}$(1260) and $\omega(782)$ resonances [871, 872].
@@
Such progress will be especially important for the baryon excitation spectrum, where many resonances couple strongly to three- or more particle channels. Lattice calculations of baryon excitations are often performed at larger pion masses to reduce the numerical cost. As sketched in the right panel of Fig. 4.16, because the nucleon mass increases with the pion mass, the $N\pi$, $N\pi\pi$, …thresholds also increase with $m_{\pi}$. Therefore, states that are resonances at the physical pion mass may eventually become bound states as $m_{\pi}$ grows. In that case, a lattice computation is straightforward and the energy levels directly correspond to the masses of the states. For example, the right panel in Fig. 4.15 shows the finite-volume energy levels for the light baryon excitation spectrum at a pion mass of about 400 MeV [800]. In these units the lowest $N\pi$ threshold corresponds to $m/m_{\Omega}\sim 1$, which is close to the lowest level in the $J^{P}=\nicefrac{{1}}{{2}}^{-}$ channel that eventually becomes the $N(1535)\nicefrac{{1}}{{2}}^{-}$. However, at the physical pion mass more and more thresholds open up, as shown in Fig. 4.16: Already the $\Delta$ baryon is close to the $N\pi\pi$ threshold, and the situation becomes progressively severe for higher-lying resonances above multiple thresholds – in these cases one would need generalizations of the Lüscher condition to four-body, five-body equations and so on.
@@
Compared to the meson sector, baryon resonances have been studied far less extensively on the lattice due to their increased computational complexity. The $\Delta$ resonance has been investigated in elastic $N\pi$ scattering [873, 874, 875], which is well justified as it decays almost exclusively into $N\pi$. The situation becomes more intricate in the negative-parity $J^{P}=\nicefrac{{1}}{{2}}^{-}$ channel [876], where several two- and also three-body decay channels are involved. This is particularly true for the Roper resonance, which is observed experimentally in both $N\pi$ and $N\pi\pi$ final states. On the lattice, the Roper has remained elusive for several reasons. First, the finite-volume energy levels in this channel vary significantly between different lattice studies, see [877, 832] for reviews and Fig. 5.9 in Sec. 5. Notably, only calculations using chiral fermions yield a low-lying nucleon excitation in the energy range of the Roper resonance [878, 877], but the discrepancies may be traced back to diagrams encoding multihadron effects [879]. Second, an elastic $N\pi$ scattering study using the Lüscher formalism [880] found no signal for the Roper resonance, which implies that three-body analyses including the $N\pi\pi$ channel are essential. Third, by combining lattice calculations with Hamiltonian effective field theory, Refs. [881, 882, 883, 884] conclude that the Roper is a dynamically generated resonance arising from coupled-channel effects between $N\pi$, $N\sigma$, $N\eta$ and $\Delta\pi$, with negligible contributions from any underlying ‘bare’ state. Taken together, these findings highlight the central role of multihadron dynamics in understanding the Roper resonance.
@@
In Sec. 5 we will contrast these findings with the electromagnetic transition form factors of the Roper measured in electroproduction, which show the characteristic zero crossing expected from a first radial excitation. Transition form factors are matrix elements of the type $\langle R|J^{\mu}|H\rangle$, where $H$ is a stable hadron like the nucleon and $R$ a resonance which can decay into two (or more) hadrons via $R\to H_{1}H_{2}\dots$. The corresponding lattice formalism [885, 886, 887, 888, 889, 890] is similar to spectroscopy calculations, i.e., one looks for pole positions in the matrix elements $\langle H_{1}H_{2}\dots|J^{\mu}|H\rangle$, which can be related to their finite-volume analogues on the lattice. Such matrix elements have been explored in the meson sector, e.g., for the process $\pi\gamma^{\ast}\to\rho\to\pi\pi$ [891, 892, 893], and it would be very interesting to extend them to the baryon sector.
@@
Concerning the $\Lambda(1405)$, most lattice calculations until recently were limited to determining the lowest finite-volume energy level using three-quark operators. Based on a form factor analysis, the authors of Ref. [894] argue that this state is primarily a $\bar{K}N$ molecule. Recently, a lattice study of coupled-channel $\pi\Sigma-\bar{K}N$ scattering at $m_{\pi}=200$ MeV has found two poles for the $\Lambda(1405)$ [895, 896], which is consistent with analyses in chiral EFT [602, 603, 604, 605].
@@
To summarize, even though lattice calculations for baryon resonances remain highly challenging, progress is being made in connecting finite-volume spectra to physical scattering amplitudes through multihadron dynamics. The Lüscher method and its extensions have stimulated a productive interplay between lattice QCD, effective field theory and amplitude analyses [897, 604]. In parallel, other formalisms to extract infinite volume scattering information from the lattice such as the HALQCD method [898, 899, 900, 901], the finite-volume Hamiltonian EFT [902, 903], the optical potential method [904], and the determination of spectral densities from Euclidean correlators [905, 906, 907] are being actively investigated. Thanks to improved algorithms, lattice calculations near or at the physical pion mass have now become feasible, and the treatment of multi-particle thresholds will be a key ingredient for a reliable determination of the baryon excitation spectrum.

### v4

@@ 4.6  Baryons in lattice QCD
Lattice QCD is a nonperturbative ab-initio method because in principle, apart from supplying the current-quark masses and setting a scale, the only information needed to compute baryon observables is the QCD action Gattringer:2010zz , Briceno:2017max . The basic idea in lattice QFT is to discretize Euclidean spacetime and consider a four-dimensional box with lattice spacing $a$, length $L$ and appropriate boundary conditions. Correlation functions inside that box are then calculated through the Euclidean version of the path integral in Eq. (4.7). The path integral over the quark and antiquark fields can be rewritten as a fermion determinant, so that the remaining integral amounts to statistically sampling over many gauge (i.e., gluon) field configurations. The typical workflow of lattice calculations consists of three steps: generating the gauge-field ensembles that enter in the path integral; ‘measuring’ the correlation functions $\langle\mathcal{O}\rangle$ on those ensembles; and extracting physical observables from the correlation functions. The last part comes with several complications mainly induced by the finite volume ($L<\infty$), the non-zero lattice spacing ($a>0$), and the difficulties in going to physical pion masses due to the accumulating statistical noise.
The starting point for extracting the baryon spectrum on the lattice is again Fig. 4.3. To begin with, suppose we start with some (gauge-dependent) correlation function that permits baryon poles, like the qqqqqq correlation function in the second row on the right. By contracting it with appropriate structures in Dirac, color and flavor space, we arrive at a (gauge-invariant) two-point correlator which is shown in the third row. This two-point function only depends on one variable z=xyz=x-y and can be interpreted as a baryon-baryon correlator $G(z)=\langle\mathcal{O}_{i}(x)\,\bar{\mathcal{O}}_{j}(y)\rangle$, where $\mathcal{O}_{i}(x)$ is some gauge-invariant combination of three quark operators. In Euclidean momentum space, this correlator depends on the total baryon momentum $P$ and inherits the poles from the original correlation function at $P^{2}=-m_{\lambda}^{2}$:
@@
where $E_{\lambda}^{2}={\mbox{$P$}}^{2}+m_{\lambda}^{2}$ and the overlap $R_{\lambda}$ is the product of the contracted Bethe-Salpeter wave functions. When taking a Fourier transform in $P_{4}$, a pole in momentum space turns into an exponential decay in Euclidean time $z_{4}=\tau$:
@@
$G(\tau,{\mbox{$P$}})\stackrel{{\scriptstyle P^{2}\to-m_{\lambda}^{2}}}{{\relbar\joinrel\relbar\joinrel\relbar\joinrel\relbar\joinrel\relbar\joinrel\relbar\joinrel\rightarrow}}R_{\lambda}\int\frac{dP_{4}}{2\pi}\frac{e^{iP_{4}z_{4}}}{P_{4}^{2}+E_{\lambda}^{2}}=\frac{R_{\lambda}}{2E_{\lambda}}e^{-E_{\lambda}\tau}\,.$ (4.37)
Setting ${\mbox{$P$}}=0$, one can then extract the mass $m_{\lambda}$ at large Euclidean times $\tau$. In general Eq. (4.36) will be a sum of poles, but for large $\tau$ the ground state with the lowest-lying mass dominates the sum.
@@
While the above equations sketch the basic ideas, a clean extraction of hadronic states from lattice QCD requires overcoming significant technical challenges. Over the years, various methods have been developed to address these issues and reduce the computational cost, including distillation HadronSpectrum:2009krc , Morningstar:2011ka , noise reduction techniques Bali:2009hu , the inclusion of large operator bases also with multihadron operators to solve generalized eigenvalue problems, and matching the irreducible representations of the cubic group to physical $J^{P}$ quantum numbers; see e.g. Dudek:2010wm , Thomas:2011rh , Dudek:2012gj , Edwards:2012fx for details. From the perspective of Fig. 4.3, the inclusion of multihadron and gluonic operators for solving generalized eigenvalue problems of two-point functions $\langle\mathcal{O}_{i}\,\bar{\mathcal{O}}_{j}\rangle$ amounts to finding a given hadron pole in a matrix of qqqqqq, $(qqq)(q\bar{q})$, $(qqq)g$ …correlation functions, which therefore also provides insight into their multiquark and gluonic admixtures. In this way, for instance, the existence of hybrid baryons with significant gluonic components has been demonstrated Dudek:2012ag .
@@
Lattice calculations of stable hadrons are well established by now, see e.g. PACS-CS:2008bkb , BMW:2008jgk , Bietenholz:2011qq , Lin:2011ti , Fodor:2012gf , Engel:2013ig , Alexandrou:2014sha , Alexandrou:2017xwd . In the absence of electroweak interactions, the ground-state octet baryons $N$, $\Lambda$, $\Sigma$, $\Xi$ and the decuplet $\Omega$ baryon are stable under the strong interaction. Thanks to improved algorithms, calculations of their masses have also become possible at physical pion masses as shown in the left panel of Fig. 4.15. This also allows one to investigate precision effects like isospin breaking and electromagnetic corrections to the masses Blum:2010ym , BMW:2014pzb , Horsley:2015eaa , NPLQCD:2020ozd . The FLAG (Flavour Lattice Averaging Group) review FlavourLatticeAveragingGroupFLAG:2021npn , FlavourLatticeAveragingGroupFLAG:2024oxs provides a comprehensive meta-analysis of results for several hadronic quantities ensuring strict quality criteria such as continuum extrapolations, finite-volume corrections and renormalization procedures.
@@
Figure 4.16: (color online) Left and center: Sketch of the Lüscher method. From the knowledge of the finite-volume spectrum at different lattice volumes one can reconstruct the phase shifts and hence the scattering amplitude. Right: Sketch of the $N\pi$, $N\pi\pi$, …thresholds (dashed) plotted over the squared pion mass. The squares are the experimental values for the nucleon, $\Delta$ and Roper mass at the physical pion mass. The solid curves are parametrizations of lattice results for $m_{N}$ and $m_{\Delta}$ (see Eichmann:2016yit for references).
@@
In contrast to stable ground states, the extraction of unstable hadrons above thresholds is substantially more difficult on the lattice. The reason for this is the finite lattice volume $L$, which does not admit branch cuts, Riemann sheets or resonance poles in scattering amplitudes Maiani:1990ca . Instead, the energy levels calculated on the lattice correspond to discrete poles on the real axis, which are the finite-volume analogues of the branch cuts in the infinite volume, i.e., the multiparticle continuum. As such, the energy levels above thresholds also cannot be directly interpreted as physical states.
@@
One way to deal with this is the Lüscher formalism Luscher:1986pf , Luscher:1990ux including its various generalizations Rummukainen:1995vs , Kim:2005gf , He:2005ey , Bernard:2008ax , Leskovec:2012gb , Fu:2011xz , Hansen:2012tf , Briceno:2012yi , Guo:2012hv , Briceno:2014oea , Agadjanov:2016mao , see Briceno:2017max , Hansen:2019nir , Mai:2021lwb for reviews. The basic idea is best explained with the help of a Bethe-Salpeter equation. Consider an elastic $2\to 2$ scattering amplitude such as $N\pi\to N\pi$ (or a coupled system of such scattering amplitudes), which creates intermediate poles for the nucleon and its excitations. Like in Eq. (4.15), we write down the generic BSE for this amplitude, once in the infinite volume and once in a finite volume (the latter with subscript $L$):
@@
Over the past decade, large progress has been made in lattice QCD calculations of hadron masses above two-particle thresholds; see Briceno:2017max , Padmanath:2018zqw , Detmold:2019ghl , Brambilla:2019esw , Bulava:2022ovd , Mai:2022eur for overviews. Most practical applications of the Lüscher method to date have focused on the meson sector. The case of elastic scattering applies to resonances with a single decay channel, such as the $\rho$ and $K^{\ast}$ which decay almost exclusively into $\pi\pi$ and $K\pi$, respectively. Other resonances like the $\sigma$ and $\kappa$ mesons Briceno:2016mjc , Guo:2018zss , Rodas:2023nec and a number of states in the heavy-meson spectrum, including exotic candidates Mohler:2012na , Prelovsek:2013cra , Mohler:2013rwa , Bicudo:2015vta , Francis:2016hui , Moir:2016srx , Bali:2017pdv , Cheung:2017tnt , Junnarkar:2018twb , Padmanath:2022cvl , have also been studied by various lattice collaborations. Since most resonances are inelastic and couple to multiple decay channels, also coupled-channel generalizations of the Lüscher formalism to inelastic scattering have been applied to a variety of systems Prelovsek:2014swa , Dudek:2014qha , Wilson:2015dqa , Dudek:2016cru , Briceno:2017qmb , JPAC:2018zyd , Woss:2020ayi , Prelovsek:2020eiw . These investigations enable the determination of resonance pole positions in the complex plane and provide insight into their dependence on the current-quark masses.
@@
The difficulty is that each additional threshold complicates the situation substantially. For unstable states above three-particle thresholds, the corresponding BSE turns into a genuine three-body problem. The generalization of the Lüscher condition to states above three-particle thresholds has been worked out by several groups Hansen:2014eka , Hansen:2015zga , Hammer:2017uqm , Hammer:2017kms , Mai:2017bge , Mai:2018djl , Jackura:2019bmu , Briceno:2019muc , Blanton:2020gha , Blanton:2020jnm , see Hansen:2019nir , Mai:2021lwb for reviews. This has found various recent applications in the meson sector, e.g., for $\pi\pi\pi$ and $KKK$ systems Blanton:2019igq , Horz:2019rrn , Blanton:2019vdk , Culver:2019vvu , Muller:2020wjo , from where one can determine the pole positions of the $a_{1}$(1260) and $\omega(782)$ resonances Mai:2021nul , Yan:2024gwp .
@@
Such progress will be especially important for the baryon excitation spectrum, where many resonances couple strongly to three- or more particle channels. Lattice calculations of baryon excitations are often performed at larger pion masses to reduce the numerical cost. As sketched in the right panel of Fig. 4.16, because the nucleon mass increases with the pion mass, the $N\pi$, $N\pi\pi$, …thresholds also increase with $m_{\pi}$. Therefore, states that are resonances at the physical pion mass may eventually become bound states as $m_{\pi}$ grows. In that case, a lattice computation is straightforward and the energy levels directly correspond to the masses of the states. For example, the right panel in Fig. 4.15 shows the finite-volume energy levels for the light baryon excitation spectrum at a pion mass of about 400 MeV Edwards:2011jj . In these units the lowest $N\pi$ threshold corresponds to $m/m_{\Omega}\sim 1$, which is close to the lowest level in the $J^{P}=\nicefrac{{1}}{{2}}^{-}$ channel that eventually becomes the $N(1535)\nicefrac{{1}}{{2}}^{-}$. However, at the physical pion mass more and more thresholds open up, as shown in Fig. 4.16: Already the $\Delta$ baryon is close to the $N\pi\pi$ threshold, and the situation becomes progressively severe for higher-lying resonances above multiple thresholds – in these cases one would need generalizations of the Lüscher condition to four-body, five-body equations and so on.
@@
Compared to the meson sector, baryon resonances have been studied far less extensively on the lattice due to their increased computational complexity. The $\Delta$ resonance has been investigated in elastic $N\pi$ scattering Andersen:2017una , Silvi:2021uya , Bulava:2022vpq , which is well justified as it decays almost exclusively into $N\pi$. The situation becomes more intricate in the negative-parity $J^{P}=\nicefrac{{1}}{{2}}^{-}$ channel Lang:2012db , where several two- and also three-body decay channels are involved. This is particularly true for the Roper resonance, which is observed experimentally in both $N\pi$ and $N\pi\pi$ final states. On the lattice, the Roper has remained elusive for several reasons. First, the finite-volume energy levels in this channel vary significantly between different lattice studies, see Liu:2016rwa , Padmanath:2018zqw for reviews and Fig. 5.9 in Sec. 5. Notably, only calculations using chiral fermions yield a low-lying nucleon excitation in the energy range of the Roper resonance Mathur:2003zf , Liu:2016rwa , but the discrepancies may be traced back to diagrams encoding multihadron effects xQCD:2019jke . Second, an elastic $N\pi$ scattering study using the Lüscher formalism Lang:2016hnn found no signal for the Roper resonance, which implies that three-body analyses including the $N\pi\pi$ channel are essential. Third, by combining lattice calculations with Hamiltonian effective field theory, Refs. Liu:2016uzk , Kiratidis:2016hda , Wu:2017qve , Owa:2025mep conclude that the Roper is a dynamically generated resonance arising from coupled-channel effects between $N\pi$, $N\sigma$, $N\eta$ and $\Delta\pi$, with negligible contributions from any underlying ‘bare’ state. Taken together, these findings highlight the central role of multihadron dynamics in understanding the Roper resonance.
@@
In Sec. 5 we will contrast these findings with the electromagnetic transition form factors of the Roper measured in electroproduction, which show the characteristic zero crossing expected from a first radial excitation. Transition form factors are matrix elements of the type $\langle R|J^{\mu}|H\rangle$, where $H$ is a stable hadron like the nucleon and $R$ a resonance which can decay into two (or more) hadrons via $R\to H_{1}H_{2}\dots$. The corresponding lattice formalism Lellouch:2000pv , Agadjanov:2014kha , Briceno:2014uqa , Briceno:2015csa , Agadjanov:2016fbd , Briceno:2021xlc is similar to spectroscopy calculations, i.e., one looks for pole positions in the matrix elements $\langle H_{1}H_{2}\dots|J^{\mu}|H\rangle$, which can be related to their finite-volume analogues on the lattice. Such matrix elements have been explored in the meson sector, e.g., for the process $\pi\gamma^{\ast}\to\rho\to\pi\pi$ Briceno:2015dca , Briceno:2016kkp , Alexandrou:2018jbt , and it would be very interesting to extend them to the baryon sector.
@@
Concerning the $\Lambda(1405)$, most lattice calculations until recently were limited to determining the lowest finite-volume energy level using three-quark operators. Based on a form factor analysis, the authors of Ref. Hall:2014uca argue that this state is primarily a $\bar{K}N$ molecule. Recently, a lattice study of coupled-channel $\pi\Sigma-\bar{K}N$ scattering at $m_{\pi}=200$ MeV has found two poles for the $\Lambda(1405)$ BaryonScatteringBaSc:2023zvt , BaryonScatteringBaSc:2023ori , which is consistent with analyses in chiral EFT Meissner:2020khl , Mai:2020ltx , Mai:2022eur , Guo:2023wes .
@@
To summarize, even though lattice calculations for baryon resonances remain highly challenging, progress is being made in connecting finite-volume spectra to physical scattering amplitudes through multihadron dynamics. The Lüscher method and its extensions have stimulated a productive interplay between lattice QCD, effective field theory and amplitude analyses JPAC:2021rxu , Mai:2022eur . In parallel, other formalisms to extract infinite volume scattering information from the lattice such as the HALQCD method Ishii:2006ec , Aoki:2009ji , Ishii:2012ssm , Aoki:2012tk , the finite-volume Hamiltonian EFT Hall:2013qba , Wu:2014vma , the optical potential method Agadjanov:2016mao , and the determination of spectral densities from Euclidean correlators Bulava:2019kbi , Bulava:2023mjc , Patella:2024cto are being actively investigated. Thanks to improved algorithms, lattice calculations near or at the physical pion mass have now become feasible, and the treatment of multi-particle thresholds will be a key ingredient for a reliable determination of the baryon excitation spectrum.